Ab initio study of a quasi van der Waals molecule: N_{2}O_{4}.

Y. ELYOUSSOUFI,\underline{J. LIEVIN} and M. HERMAN, Laboratoire de Chimie Physique Moléculaire, Université Libre de Bruxelles, CPI 160/09, B-1050 Bruxelles, BELGIUM.

A critical review of the ab initio literature on the ground electronic state of N_{2}O_{4} is presented. It is completed by new calculations including higher order electron correlation effects using the MP4 and internally-contracted MRCI methods and some open questions are revisited. They concern the harmonic and anharmonic force fields, the torsional energy barrier, the dissociation path into two NO_{2} radicals and the assignment of the IR spectrum. Most properties affected by the quasi van der Waals nature of the NN bond are shown to require a high level of theory (MRCI/cc-pVTZ). Reliable harmonic vibrational frequencies (within 3%) are however obtained from the MP4/6-31G* method. This level has been adopted to calculate the IR spectrum from anharmonic variational calculations within a 4 dimensional effective model. An unambiguous assignment of the experimental spectrum^{a} below 600 cm^{−1} is proposed.

aJ. Koput, J.W.G. Seibert and B.P. Winnewisser, Chem. Phys. Lett., 204, 183 (1993).